Sheet of Quartz Due in 9 hours, 42 minutes time t takes the light to travel through the two sheets, how far (in cm) would it have traveled in a vacuum?
Answer #1

The refractive index (n) is the ratio c/s where s is the speed of light in a medium and c is the speed of light in vacuum.

Since distance (d) is proportional to speed we can write (for ice)

c/si = 1.309 = dv/2.01 where subscripts v & i refer to vacuum and ice respectively.

dv = 2.63cm this is the distance the light would travel in vacuum in the same time it would take to travel in the 2.01 cm ice. (optical path).

similarly, for quartz, dv = 1.37\times1.544= 2.11cm.

total distance in vacuum = 2.63+2.11= 4.74 cm.